OpenStudy (anonymous):

Given sin A=2/3 and cos B = -1/2 where A is in quadrant 2 and B is in Quadrant 3 find Sin(A+B), cos (A-B), and tan(A-B)

OpenStudy (anonymous):

\[\sin(A+B)=\sin(A)\cos(B)+\sin(B)\cos(A)\]plug in the numbers

OpenStudy (anonymous):

you know \[\sin(A)=\frac{2}{3}\] do you know how to find \[\cos(A)\]?

OpenStudy (anonymous):

I don't, I miss class the day they went over this!!

OpenStudy (anonymous):

draw a right triangle with an angle \(A\) make the opposite side of the angle 2, and the hypotenuse 3

OpenStudy (anonymous):



OpenStudy (anonymous):

use pythagoras to find the missing sides let me know what you get

OpenStudy (anonymous):

if it is not clear let me know an i can show you